Vocaloid:

if L is the length of each string, what is the maximum height of ball E in trial 1, above the lowest position, in terms of L? (pic below)

Vocaloid:

https://imgur.com/a/UKTxDtA

imqwerty:

considering no loss of energy and elastic collision the Ball E should rise up to the level from which ball A is released

imqwerty:

the momentum that Ball A transfers to ball B is all in horizontal direction so all the energy transferred is converted into kinetic 

imqwerty:

had the momentum transfer been in the direction of the arrow show above there'd be some energy loss because there is no motion possible in the y direction unless the string is elastic but in our case the momentum transfer is purely in the x direction so theres no energy loss

Vocaloid:

how would I solve for h in terms of L and theta?

imqwerty:

you're gonna have to use trigonometry for that

Vocaloid:

 have something like this set up but forgot how to proceed from here

imqwerty:

you can use \(\cos(\theta)\) to find h

Vocaloid:

oh duh now I see it cos(theta) = (L-h)/L Lcos(theta) = L - h h = L - Lcos(theta)

Vocaloid:

thanks

imqwerty:

np